OpenStudy (pulsified333):

Let f(1)=−3 and f′(1)=8. Then the equation of the tangent line to the graph of y=f(x) at x=1 is y=

OpenStudy (pulsified333):

@dan815

OpenStudy (pulsified333):

well i Know we have one point (1,-3)

OpenStudy (pulsified333):

okay but what is the other point?

ganeshie8 (ganeshie8):

what else do you need to write the equation of tangent line in "point - slope" form ?

ganeshie8 (ganeshie8):

remember the equation of a lien in "point-slope" form ?

OpenStudy (pulsified333):

y+3=m(1-x)

ganeshie8 (ganeshie8):

that is close, try again : 

OpenStudy (pulsified333):

whoops y+3=m(x-1)

ganeshie8 (ganeshie8):

Yes, how to find the value of the slope, m ?

ganeshie8 (ganeshie8):

derivative gives you the slope

OpenStudy (pulsified333):

but we know that the derivative of 1 is 8

ganeshie8 (ganeshie8):

derivative of a function at a point gives the slope of the tangent line to the graph at that point

OpenStudy (pulsified333):

so is the slope 8

ganeshie8 (ganeshie8):

Yes

ganeshie8 (ganeshie8):

simply plugin m=8 in the tangent line eqn that you have got earlier

OpenStudy (pulsified333):

so the equation is y=8x-5

ganeshie8 (ganeshie8):

step by step, please y+3=m(x-1) plugin m = 8, what do you get ?

OpenStudy (pulsified333):

y+3=8(x-1)

ganeshie8 (ganeshie8):

you should be able to do the simplification on your own

ganeshie8 (ganeshie8):

just work it step by step

OpenStudy (pulsified333):

whoops i got it silly mistake y=8x-11

ganeshie8 (ganeshie8):

looks good

OpenStudy (pulsified333):

thanks man

ganeshie8 (ganeshie8):

Np